A 4.2 g lead bullet moving at 275 m/s strikes a steel plate and stops. If all the kinetic energy is converted to thermal energy and none leaves the bullet, what is its temperature change?
1
Expert's answer
2020-10-29T08:20:05-0400

π‘š = 4.2 𝑔 = 0.00420 π‘˜π‘”

𝑣𝑖 = 275 π‘šβ„π‘ 

𝑣𝑓 = 0

𝑐 = 130 π½β„π‘˜π‘” βˆ™ ℃

βˆ†π‘‡ =?

All of the kinetic energy of the bullet is converted into thermal energy, βΈ«

βˆ†πΎπΈ = mcβˆ†T

βˆ†T = "\frac{∆KE}{mc}"


= - "\frac{(1/2)m(v_f^2-v_i^2)}{mc}" = - "\frac{(1/2)(v_f^2-v_i^2)}{c}"


= - "\frac{(1/2)(0-275m/s^2)}{130J/Kg. °C}"


βˆ†T = 290Β°C